From mar@roe.ac.uk Mon Dec 5 10:52:02 2016 Date: Mon, 5 Dec 2016 10:51:55 +0000 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , dmr@roe.ac.uk, Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Alastair Edge , Andrew Lawrence , Bob Mann , Dante Minniti , James Dunlop , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Maria-Rosa Cioni , matt.jarvis@astro.ox.ac.uk, Mike Irwin , Nigel Metcalfe , Norman Gray , Phil Lucas , Richard McMahon at IOA , Simon Dye , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives progress meeting minutes, 25 November 2016 WFAU VDFS Science Archives progress meeting minutes, 25 November 2016 Present: NJC, RPB, ETWS, MAR, MSH, RGM Apologies: STV, DMR, KTN, AL, RSC, ACD DoNM: 10:00 am Friday 9th November 2016, VILLAS MEETING ROOM 2 ACD in the chair (room is booked). Actions discharged this time: ---------------------------- ACTION: MSH to check with DMR about progress towards, and anticipated release date of, a usable TAP service. Done, new TAP service hopefully available January 2017. ACTION NJC: investigate and fix the problem with Wserv8. NJC had contacted CASU to see if they could spot anything in the imcore code that was causing the NaNs as the inputs looked OK. Done, an update of the CASU code had fixed the issue and WSERV8 had been released. ACTION MAR: reply to DMR's e-mail about publishing statistics for columns. Done, MAR had discussed with DMR last week, all seems sensible as we wouldn't have provide every column/stat, just the ones we think might be useful. Actions carried forward from 11/11/2016 meeting: --------------------------------------------- ACTION: RGM to investigate the size and content (how does it differ from DR9?) of SDSS DR10 to see if it is worth us hosting and whether it is necessary to keep all of the earlier releases. Continues. The first part had been done: the difference between DR9 and DR10 is largely the presence of additional spectra. The part of this action that continues is that RGM had not yet established which early releases could safely be scrapped. RGM also noted that that DR11 and DR12 were now available and now need to be looked at. For further details see the WFAU SDSS strategy wiki page: http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan DR13 is installed but not yet cross-mounted. However, we still need to decide which SDSS releases to keep so the present action should remain. ACTION: MAR to write some H-ATLAS documentation once H-ATLAS is available. Continues. ACTION: RPB to check the log files and inform NJC of the problems with the previously listed non-survey releases. Continues. ACTION: NJC to investigate and fix the problems with the previously listed non-survey releases. Continues. This action cannot progress until RPB checks the log files and passes on details of the problems (action above). ACTION: RPB to check with CASU whether they had any WFCAM observations made during the gap in the Edinburgh tapes. It was suggested we might like to check here first! Continues. ACTION: NJC to complete the deep-stacking software for VVV. Continues. ACTION: ETWS to ask for and re-transfer the missing SDSS databases. Continues, ETWS had asked. ACTION: MSH to investigate speed of disk44 Continues. ACTION: ACD to set-up a WIKI page on MS SQL Server on Linux and all to add their input and what and how it needs to be tested for our uses. Continues Specific points and new actions: -------------------------------- Project management: * RGM noted that he would create a doodle poll for PDR time slots. * Christmas lunch is Monday 19th, The Sheep Heid, at 1pm, hooray. * WFAU quarterly is Tuesday 20th, M1 meeting room, 1:30pm, boo. WFCAM, VISTA and VST updates: * Nothing to report. Networking: * MSH noted that the site router was being upgrading and that this would require us to upgrade our switch with new transceivers. No downtime was anticipated. WSA/VSA/OSA/GES Operations: * ETWS reported that he had: - OSA VPHAS+ is still going fine, reseaming is 40% complete. - Transferred the latest released VISTA data, April to June 2016. - Created new browser pages for WSA and SSA. - Ingested the OSA catalogue data (fixed and uncorrected) for the April to June 2016 data. objID and illumCorr updates are running. RSC had fixed a bug in the supplied iDR5 linelist. Hardware and Systems: * RPB noted that he had fixed up the "dead" database fallout from the failure of ramses7. Review of Software Development Deadlines: * The list of software deadlines was briefly discussed. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/PrepareSoft Would be done offline Other Software: * ETWS noted: - Installed latest (v1.0.34) casutools. - Minor fixes. NJC noted he'd done further work on the list driven pipeline, now running it on large batches of data and optimising it. Review of Survey Release Deadlines: * Progress towards forthcoming survey deadlines was reviewed. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Survey Data Release: * VHS DR4 was released/made public. Non-survey Data Release: * WSERV8 was released Astrogrid deployment, VO & Data Analysis services: Nothing for the minutes. Miscellaneous: DEX XIII will be held at ROE on Monday 9th - Tuesday 10th January 2017.